Part C Find an expression for the amount of work the gas does in terms of quanities given in the problem rs moles of an ideal monatomic gas expand at a constant pressure P: the volume of the gas changes from V to Correct Part D Find an for the amount of heat added to the gas in terms of quansties given in the problem
Answer #1

d)

H = n Cp ( T2 - T1)

H = n (5/2) R * PV ( 7/5 - 1) / nR

H = PV

========

e)

H = U + P ( 7/5 - 1) V

PV = U + PV( 2/5)

U = (3/5) PV

========